OpenStudy (anonymous):

f(m)=4/m-4 g(m)=-4/3m-4 Find:(f-g)(m)

OpenStudy (2bornot2b):

\[(f-g)(m)=f(m)-g(m)=\frac{4}{m}-4+\frac{4}{3m}+4\] Now I think you can do the rest

OpenStudy (2bornot2b):

\[\frac{12\cancel{-12m}+4\cancel{+12m}}{3m}\implies16/3m\]

OpenStudy (anonymous):

nah it was wrong :/